OpenStudy (anonymous):

how do you factor (3x squared y squared -75) ?

OpenStudy (anonymous):

factor out a 3 and then factor the difference of two squares

OpenStudy (anonymous):

what answer did you get?

OpenStudy (anonymous):

by which i mean start with \(3(x^2y^2-25)\) and then use \(a^2-b^2=(a+b)(a-b)\) with \(a=xy, b=5\)